Operations Report – December 29th

Crew 236 Operations Report 29-DEC-2021

SOL: 10

Name of person filing report: Pavithra Ravi

Non-nominal systems: None

ROVERS

Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: Left unplugged after last EVA to save power. To be
charged by Shannon.

Opportunity rover used: Yes
Hours: 77.8
Beginning charge: 100%
Ending charge: 89%
Currently charging: Left unplugged after EVA to save power. To be
charged by Shannon.

Curiosity rover used: Yes
Hours: 157.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Left unplugged after EVA to save power. To be
charged by Shannon.

Perseverance rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: To be charged by Shannon.

General notes and comments: None

Summary of Hab operations: The toilet flush is not functioning and
there is about an inch of waste stuck to the bottom of the tank that
we have not been able to flush away. In the past 24 hours, we have
flushed the toilet about 4 times (by filling buckets in the shower and
emptying it into the toilet), to no avail. Flushing the toilet for
individual use with a little bucket has not been too rough, so we can
make it work till the end of our rotation, after which a new toilet
will be installed.

We were in power conservation mode today due to cloud cover limiting
the available solar power.

WATER USE:

Water (static tank): approx. 100-150 gallons after all that flushing!
Water (loft tank): 21 gallons
Water Meter: 01542759 units

Static to Loft Pump used – yes or no: Yes. It was filled from 32 to 55
gallons this afternoon.
Static tank pipe heater (on or off): On
Static tank heater (on or off) Off

Toilet tank emptied (no or yes): yes, around 14:30

Summary of internet: Nominal

Summary of suits and radios: Nominal

We ran out of one can of plexiglass cleaner and are onto a second one.

Summary of GreenHab operations: Nominal

WATER USE: 7 gallons were used today. This leaves about 158 gallons
in the tank.

Heater (on or off): On

Supplemental light (on or off): On

Harvest: 5g of parsley, 2g of wheatgrass, 1g Thai hot pepper

Summary of ScienceDome operations: Nominal

Dual split, Heat or AC (on or off): off

Summary of RAM operations: Tools used, work done

Nothing used.

Summary of any observatory issues: Nothing new to report.

Summary of health and safety issues: Nominal

Questions, concerns, and requests to Mission Support: None